At the beginning of school year the average reading comprehension of freshmen college students was recorded at 79.40. After 3 months of studying the same instruments for reading comprehension test was used for 56 students and scored an average of 86.48 with a standard deviation of 4.65. Does staying in the college improve reading comprehension of the freshmen students? Use 0.05 level of significance
1
Expert's answer
2020-05-25T16:31:46-0400

"a_0=79.4\\
n=56\\
\overline{x}=86.48\\
s=4.65\\
\alpha=0.05\\
H_0: \mu=a_0=79.4, H_1:\mu>a_0=79.4\\
\mu\text{ --- reading comprehension of the students after}\\
\text{3 months of studying}.\\
\text{We assume that reading comprehension of the students}\\
\text{has normal distribution}.\\
\text{We will use the following random variable as a criterion:}\\
T=\frac{(\overline{X}-a_0)\sqrt{n}}{s}\\
T\text{ has t-distribution with } k=n-1 \text{ degrees of freedom}.\\
\text{Observed value:}\\
t_{obs}=\frac{(86.48-79.4)\sqrt{56}}{4.65}\approx 11.4\\
\text{Critical value (one-sided)}:\\
t_{cr}=t_{cr}(\alpha;k)=t_{cr}(0.05;55)\approx 1.673\\
(1.673,\infty)\text{ --- critical region}\\
t_{obs}\text{ falls into the critical region. So we reject } H_0.\\
\text{Staying in the college improves reading comprehension}\\
\text{of the freshmen students}."